OpenStudy (anonymous):

Solve sqrt-4/(5+2i)-(3-4i)

OpenStudy (anonymous):

\[\sqrt{\frac{-4}{(5+2i)-(3-4i)}}\]\[=\sqrt{\frac{-4}{2+5i}}=\sqrt{\frac{-4}{2+5i} \times \frac{2-5i}{2-5i} }\] \[=\sqrt{\frac{-4(2-5i)}{29}}=\sqrt{\frac {8}{29}}i+\sqrt {\frac {20i}{29}}= \sqrt{\frac {8}{29}}i+\sqrt {\frac {20}{29}}\sqrt{i}\] something like that ... I dont know how to deal with sqrt i . if any one do .. then i be grateful . regards